OpenStudy (nathanjhw):

The function f (x) has the value f (−1) = 1. The slope of the curve y = f (x) at any point is given by the expression dy/dx = (2x + 1)( y + 1)

OpenStudy (nathanjhw):

A. Write an equation for the line tangent to the curve y = f (x) at x = − 1. B. Use the tangent line from part A to estimate f (−0.9) C. Use separation of variables to find an explicit or implicit formula for y = f (x), with no integrals remaining. D. Find lim f (x) as x approaches infinity.

OpenStudy (nathanjhw):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

we're told that f(-1) = 1 so that means (-1,1) lies on the graph of f(x)

jimthompson5910 (jim_thompson5910):

plug in x = -1 and y = 1 into the dy/dx formula and that will determine the slope m

OpenStudy (nathanjhw):

-1?

OpenStudy (nathanjhw):

-2 sorry

jimthompson5910 (jim_thompson5910):

nope on either one

OpenStudy (nathanjhw):

How do I do it then?

jimthompson5910 (jim_thompson5910):

you plugged it into dy/dx = (2x + 1)( y + 1) right?

OpenStudy (nathanjhw):

(2(-1)+1)(1+1)

jimthompson5910 (jim_thompson5910):

oh wait, I divided on accident for some reason

jimthompson5910 (jim_thompson5910):

idk why I did that

jimthompson5910 (jim_thompson5910):

yeah you're right, m = -2

jimthompson5910 (jim_thompson5910):

anyways, we know the following x = -1 y = 1 m = -2 now plug those into y = mx+b and solve for b

OpenStudy (nathanjhw):

b= -1

jimthompson5910 (jim_thompson5910):

so the tangent line is y = -2x - 1

OpenStudy (nathanjhw):

Alright, now how do we do B?

jimthompson5910 (jim_thompson5910):

plug in x = -0.9 into the tangent line equation to get the approximate value for f(-0.9)

OpenStudy (nathanjhw):

-2.8

jimthompson5910 (jim_thompson5910):

the idea is that if we plug in values close to x = -1, then we'll get rough approximations to f(x) be warned: the further you stray from x = -1, the larger the error will be until it will be so large that it's not worth it

jimthompson5910 (jim_thompson5910):

-2.8? you sure?

OpenStudy (nathanjhw):

-2(0.9)-1

jimthompson5910 (jim_thompson5910):

negative 0.9

OpenStudy (nathanjhw):

oops

OpenStudy (nathanjhw):

0.8

jimthompson5910 (jim_thompson5910):

that's the approx value of f(-0.9)

OpenStudy (nathanjhw):

Now how do we do part C?

jimthompson5910 (jim_thompson5910):

\[\Large \frac{dy}{dx} = (2x+1)(y+1)\] \[\Large dy = (2x+1)(y+1)dx\] \[\Large \frac{dy}{y+1} = (2x+1)dx\] \[\Large \int\frac{dy}{y+1} = \int(2x+1)dx\] \[\Large \ln(|y+1|) = x^2+x+C\] I'll let you finish up. The goal is to isolate y (which is the same as finding f(x))

OpenStudy (nathanjhw):

Is it y = e^(C+x^2+x)-1

jimthompson5910 (jim_thompson5910):

\[\Large \ln(|y+1|) = x^2+x+C\] \[\Large |y+1| = e^{x^2+x+C}\] \[\Large |y+1| = e^C*e^{x^2+x}\] Let's find the value of e^C using x = -1 and y = 1 \[\Large |y+1| = e^C*e^{x^2+x}\] \[\Large |1+1| = e^C*e^{(-1)^2+(-1)}\] \[\Large |2| = e^C*e^{0}\] \[\Large 2 = e^C*1\] \[\Large 2 = e^C\] \[\Large e^C = 2\] ------------------------------------------------------- So \[\Large |y+1| = e^C*e^{x^2+x}\] turns into \[\Large |y+1| = 2*e^{x^2+x}\]

jimthompson5910 (jim_thompson5910):

The absolute value is often a pain, but the good news is that because the right side is never negative, we can drop the absolute values. So solving for y gives \[\Large |y+1| = 2*e^{x^2+x}\] \[\Large y+1 = 2*e^{x^2+x}\] \[\Large y = 2*e^{x^2+x}-1\] \[\Large f(x) = 2*e^{x^2+x}-1\]

OpenStudy (nathanjhw):

Alright, sorry I had to go somewhere @jim_thompson5910 back now though.

OpenStudy (nathanjhw):

For the last one, f(x) is the answer to c right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yeah hopefully you see how it all leads up to that

OpenStudy (nathanjhw):

I got infinity

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (nathanjhw):

alright thank you so much for all your help @jim_thompson5910 really appreciate it!

jimthompson5910 (jim_thompson5910):

np